OpenStudy (anonymous):

does anyone have a good thesis statement for staying sober in todays society

OpenStudy (anonymous):

5(5) + 63 = 25+63= 88

OpenStudy (anonymous):

2(36) divided by 2 = 36

OpenStudy (anonymous):

What are you planning to talk about in the essay

OpenStudy (anonymous):

Because remember the thesis tell the reader what you will write about

OpenStudy (anonymous):

staying sober in today's society. When someone is a recovering alcoholic, it's hard for them trying to stay sober when there is so much of it all around everywhere one looks.